# Online-BLS: An Accurate and Efficient Online Broad Learning System for Data Stream Classification

### v1

@@ B.1 Stationary Datasets
USPS [^3] https://datasets.activeloop.ai/docs/ml/datasets/usps-dataset/: The USPS dataset is used for the task of handwritten digit classification. The USPS contains $9,298$ samples from $10$ classes ranging from 00 to $9$. Each sample is a $16\times 16$ grayscale image. To facilitate BLS modeling, we flatten each sample into a one-dimensional vector.
Letter [^4] https://archive.ics.uci.edu/dataset/59/letter+recognition: The Letter dataset is a capital letter in the English alphabet classification dataset. The Letter dataset contains $20,000$ samples. Each sample contains $16$ primary numeric attributes (statistical moments and edge counts), which are then scaled to an integer value range of 00 to $15$.
@@
MNIST [^7] http://yann.lecun.com/exdb/mnist/: The MNIST dataset is a handwritten digit recognition benchmark. Each sample is a grayscale image with $28\times 28$ pixels. The training and test sets are merged to yield a dataset containing $70,000$ samples uniformly distributed from 0 to 9. To adapt to our model, each sample is flattened into a vector and each attribute is normalized to be between 00 and $1$.
@@ B.2 Non-stationary Datasets
Hyperplane: The Hyperplane dataset is an artificial binary classification task. Its task is to divide points in a $d$-dimensional space into two parts by a $d-1$ dimensional hyperplane. A sample is labelled positive if it satisfies $\sum^{d}_{i=1}w_{i}x_{i}>w_{0}$ and negative if the opposite is true. By smoothly changing the parameters of the classification hyperplane, concept drift can be added to the generated data stream. Specifically, following Su et al. [2024], we generate a data stream containing $20$ features. Its noise level and drift magnitude are set to $1\%$ and $0.5\%$, respectively.
SEA: The SEA dataset is also an artificial binary classification dataset. Each sample contains three features, of which only the first two are relevant. If the sum of the first two features exceeds a certain threshold, then it is a positive example, otherwise it is a negative example. There are $4$ thresholds to choose from. Concept drift is introduced by switching thresholds at the $25,000$-th, $50,000$-th, and $75,000$-th data points. Moreover, following Su et al. [2024], the noise level we introduce in the process of data stream generation is $10\%$.
Electricity [^9] https://sourceforge.net/projects/moa-datastream/files/Datasets/Classification/elecNormNew.arff.zip/download: The Electricity dataset is used as a binary classification task, where the goal is to predict if the electricity price of Australian New South Wales will go up or down. In this market, prices are not fixed and are affected by demand and supply of the market. They are set every five minutes. Each sample originally had $8$ attributes. Following Su et al. [2024], we delete the attributes of date and time located in the first two columns and keep the remaining $6$ attributes. This dataset contains $45,312$ instances and its concepts fluctuate over time.
@@ C.6 Matthews Correlation Coefficient
The matthews correlation coefficient is generally considered a better metric than F1 score and accuracy. Its essence is a correlation coefficient value between $-1$ and $1$. The coefficient $+1$, 00, $-1$ represent a perfect prediction, an average random prediction, and an inverse prediction, respectively. Taking binary classification as an example, its formula is
@@
where $TP$, $TN$, $FP$, and $FN$ denote the number of true positive, the number of ture negative, the number of false positive, and the number of false negative, respectively. In fact, we call sklearn’s function [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.matthews_corrcoef.html to acquire the matthews correlation coefficient.

### v2

@@ B.1 Stationary Datasets
USPS [^3] https://datasets.activeloop.ai/docs/ml/datasets/usps-dataset/: The USPS dataset is used for the task of handwritten digit classification. The USPS contains $9,298$ samples from $10$ classes ranging from 0 to $9$. Each sample is a $16\times 16$ grayscale image. To facilitate BLS modeling, we flatten each sample into a one-dimensional vector.
Letter [^4] https://archive.ics.uci.edu/dataset/59/letter+recognition: The Letter dataset is a capital letter in the English alphabet classification dataset. The Letter dataset contains $20,000$ samples. Each sample contains $16$ primary numeric attributes (statistical moments and edge counts), which are then scaled to an integer value range of 0 to $15$.
@@
MNIST [^7] http://yann.lecun.com/exdb/mnist/: The MNIST dataset is a handwritten digit recognition benchmark. Each sample is a grayscale image with $28\times 28$ pixels. The training and test sets are merged to yield a dataset containing $70,000$ samples uniformly distributed from 0 to 9. To adapt to our model, each sample is flattened into a vector and each attribute is normalized to be between 0 and $1$.
@@ B.2 Non-stationary Datasets
Hyperplane: The Hyperplane dataset is an artificial binary classification task. Its task is to divide points in a $d$-dimensional space into two parts by a d1d-1 dimensional hyperplane. A sample is labelled positive if it satisfies $\sum^{d}_{i=1}w_{i}x_{i}>w_{0}$ and negative if the opposite is true. By smoothly changing the parameters of the classification hyperplane, concept drift can be added to the generated data stream. Specifically, following su2024elastic, we generate a data stream containing $20$ features. Its noise level and drift magnitude are set to $1\%$ and $0.5\%$, respectively.
SEA: The SEA dataset is also an artificial binary classification dataset. Each sample contains three features, of which only the first two are relevant. If the sum of the first two features exceeds a certain threshold, then it is a positive example, otherwise it is a negative example. There are $4$ thresholds to choose from. Concept drift is introduced by switching thresholds at the $25,000$-th, $50,000$-th, and $75,000$-th data points. Moreover, following su2024elastic, the noise level we introduce in the process of data stream generation is $10\%$.
Electricity [^9] https://sourceforge.net/projects/moa-datastream/files/Datasets/Classification/elecNormNew.arff.zip/download: The Electricity dataset is used as a binary classification task, where the goal is to predict if the electricity price of Australian New South Wales will go up or down. In this market, prices are not fixed and are affected by demand and supply of the market. They are set every five minutes. Each sample originally had $8$ attributes. Following su2024elastic, we delete the attributes of date and time located in the first two columns and keep the remaining $6$ attributes. This dataset contains $45,312$ instances and its concepts fluctuate over time.
@@ C.6 Matthews Correlation Coefficient
The matthews correlation coefficient is generally considered a better metric than F1 score and accuracy. Its essence is a correlation coefficient value between 1-1 and $1$. The coefficient +1+1, 0, 1-1 represent a perfect prediction, an average random prediction, and an inverse prediction, respectively. Taking binary classification as an example, its formula is
@@
where $TP$, TNTN, FPFP, and $FN$ denote the number of true positive, the number of ture negative, the number of false positive, and the number of false negative, respectively. In fact, we call sklearn’s function [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.matthews_corrcoef.html to acquire the matthews correlation coefficient.